Ten different video games showing violence were observed. The duration times of violenceviolence ​(in seconds) were recorded. When using this sample for a t test of the claim that the population mean is greater than 74 ​sec, what does df​ denote, and what is its​ value? What does df​ denote?
Answer #1

Ans: df denotes degrees of freedom. Its value is n-1

where n is sample size.

since 10 different video games were observed. Therefore n =10 in this case

which makes our df = n - 1 = 10 - 1 = 9